# VQAThinker: Exploring Generalizable and Explainable Video Quality Assessment via Reinforcement Learning

### v1

@@ Introduction
Video quality assessment (VQA) (Min et al. 2024) plays a fundamental role in video processing systems, serving as the primary criterion for evaluating perceptual quality across stages such as acquisition, compression, and enhancement. Most established VQA models (e.g., VMAF (Li et al. 2018)) rely on full-reference (FR) settings, where a pristine reference video is available for comparison. In many real-world scenarios, particularly those involving user-generated content (UGC), reference videos are unavailable, making FR methods inapplicable. Therefore, no-reference (NR) VQA, which assesses quality without relying on reference videos, offers broader real-world applicability and is the focus of this work  [^2] In this paper, we use VQA to refer to NR-VQA for simplicity..
@@
NR-VQA has witnessed significant advancements, progressing from early handcrafted feature-based methods, such as natural scene statistics (NSS) (Mittal, Moorthy, and Bovik 2012) to deep neural network (DNN) architectures, including convolutional neural networks (CNNs) (Li, Jiang, and Jiang 2019), Transformers (Wu et al. 2022), and more recently, large multimodal models (LMMs) (Ge et al. 2024). However, most existing methods rely on supervised fine-tuning (SFT) using specific video content or distortion datasets (Ying et al. 2021; Li et al. 2019), resulting in models that act as specialists for narrow scenarios but exhibit poor generalization to unseen video domains. To enhance generalization, recent studies have explored strategies, such as multi-dataset joint training (Li, Jiang, and Jiang 2019), unsupervised and self-supervised learning (Cao et al. 2025), etc., aiming to improve training diversity while minimizing labeling costs. Nevertheless, these approaches often incur significantly higher training costs due to the increased data scale and model complexity.
@@
In addition, interpretability has become a crucial capability for practical NR-VQA applications. While conventional VQA models typically output single or multi-dimensional quality scores, real-world scenarios often require diagnostic assessments that reveal the underlying causes of quality degradation, such as compression artifacts, temporal instability, etc.. Such interpretability not only enhances the reliability of VQA predictions but also enables closed-loop video processing by guiding targeted restoration strategies based on identified distortions. Recently, the rise of LMMs has advanced the development of explainable VQA, where models are typically fine-tuned on human-annotated quality instruction datasets, such as VQA2 (Jia et al. 2024) and OmniVQA-Chat-400K (Jia et al. 2025). However, these datasets are typically bootstrapped from LMM-generated descriptions and later refined by human annotators, making the final annotations susceptible to the biases and limitations of the source models. Furthermore, they often lack associated quality scores, hindering the joint learning of quality understanding and scoring, a process believed to mirror human perceptual decision-making (Mazurek et al. 2003).
@@
Developing generalizable VQA models remains a long-standing objective. Early studies primarily utilize handcrafted features based on perceptual quality priors, such as NSS (Mittal, Moorthy, and Bovik 2012), motion vectors (Konrad and Dubois 1992), optical flow (Beauchemin and Barron 1995), etc. While effective on synthetic distortions, these methods demonstrated limited efficacy on videos containing complex real-world distortions. Subsequently, DNN-based methods have been widely adopted for VQA, which typically comprise two main modules: a feature extraction module and a regression module. The feature extraction module often employs off-the-shelf networks, such as CNNs (Li, Jiang, and Jiang 2019), Transformers (Wu et al. 2022), and LMMs (Wu et al. 2023d), which are either pre-trained on large-scale datasets (e.g., IQA datasets (Li et al. 2022)) or fine-tuned directly on the target VQA datasets (Sun et al. 2022; Wu et al. 2022) to capture the spatial and temporal quality representation. The regression module then maps the extracted features to quality scores, using either MLPs that aggregate frame-level predictions (Sun et al. 2024) or sequence modeling methods, such as GRUs (Li, Jiang, and Jiang 2019), InceptionTime (Ying et al. 2021), or Transformers (Wu et al. 2023a), to capture temporal dependencies across frames. DNN-based methods significantly improve the performance on in-the-wild distortions, but they still suffer from limited generalization to unseen video content and novel distortion types (Cao et al. 2025).
To improve generalization, recent studies have explored various training strategies beyond standard supervised learning. These include mixed-dataset training with scale alignment (Li, Jiang, and Jiang 2019), contrastive learning with proxy tasks (Chen et al. 2021; Madhusudana et al. 2023), distortion-aware reconstruction via encoder-decoder architectures (Xie et al. 2024), and learning-to-rank frameworks using pseudo-labeled video pairs (Cao et al. 2025). Although these methods improve generalization to some extent by increasing the diversity of training data, they typically incur substantial costs in constructing large-scale datasets and demand significantly higher training overhead. In this paper, we demonstrate that VQAThinker enhances generalization without increasing the number of training samples, highlighting its efficiency.
@@
Q-Insight arXiv’25 LSVQ (28K) 0.644 0.639 0.601 0.648 0.751 0.753 0.624 0.708 0.560 0.591 0.635 0.651
VisualQuality-R1 arXiv’25 LSVQ (28K) 0.795 0.796 0.716 0.744 0.784 0.792 0.732 0.781 0.717 0.730 0.765 0.776
@@
Q-Insight arXiv’25 LSVQ (28K) 0.310 0.326 0.372 0.384 0.256 0.268 0.218 0.206 0.547 0.564 0.334 0.340
VisualQuality-R1 arXiv’25 LSVQ (28K) 0.472 0.548 0.493 0.574 0.340 0.347 0.227 0.298 0.622 0.637 0.408 0.458
@@
Table 1: Performance comparison of our model against other competitive methods. For the training data, VQA2-ID refers to VQA2 Instruction Dataset (Jia et al. 2024), which also includes LSVQ (Ying et al. 2021). “fused” denotes a combination of datasets, including KonIQ (Hosu et al. 2020), SPAQ (Fang et al. 2020), KADID (Lin, Hosu, and Saupe 2019), LSVQ (Ying et al. 2021), and AVA (Murray, Marchesotti, and Perronnin 2012). The best results are highlight in boldface, the second best is underlined. NA in the table indicates unavailable results. “Overall” represents the weighted average results based on the number of videos in each dataset.
@@
Bell-shaped Regression Reward. VQA is naturally formulated as a regression task, However, existing regression-based rewards struggle to provide fine-grained training signals. Some methods use constant binary rewards (Li et al. 2025a), assigning fixed values when predictions fall within a threshold, without capturing their proximity to the ground truth. Others adopt L1-norm-based rewards (Zhang et al. 2025a), where the reward scales linearly with prediction error at a constant rate. In practice, learning becomes increasingly difficult as the predicted score approaches the ground truth. Therefore, the reward function should provide stronger and more rapidly changing signals in this fine-grained regime to better guide the model toward high-precision predictions.
@@
For video quality scoring, we evaluate performance on ten VQA benchmarks: LSVQ Test (Ying et al. 2021), LSVQ 1080p (Ying et al. 2021), KoNViD-1k (Hosu et al. 2017), LIVE-VQC (Sinno and Bovik 2018), YouTube-UGC (Wang, Inguva, and Adsumilli 2019), LIVE-YT-Gaming (Yu et al. 2022), CGVDS (Saha et al. 2023), LIVE-YT-HFR (Madhusudana et al. 2021), Waterloo-IVC-4K (Li et al. 2019), and VDPVE (Gao et al. 2023). The first five datasets are considered in-domain, as they primarily consist of UGC videos with in-the-wild distortions. The remaining five datasets are treated as OOD due to their distinct content types (e.g., gaming videos) or distortion patterns (e.g., compression and enhancement artifacts). For video quality understanding, we adopt two evaluation benchmarks: FineVD (Duan et al. 2025) and Q-Bench-Video (Zhang et al. 2025b). FineVD is used to assess the model’s ability to identify quality-related distortion attributes, while Q-Bench-Video measures the accuracy of video quality descriptions.
@@
Competing Methods We compare VQAThinker with three categories of methods: 1) unsupervised methods: NIQE (Mittal, Soundararajan, and Bovik 2012), VIIDEO (Mittal, Saad, and Bovik 2015), STEM (Kancharla and Channappayya 2021), and CLIP-IQA (Wang, Chan, and Loy 2023); 2) supervised methods: SimpleVQA (Sun et al. 2022), FAST-VQA (Wu et al. 2022), DOVER (Wu et al. 2023b), MinimalisticVQA (IX) (Sun et al. 2024), Q-Align (Wu et al. 2023d) and VQA2-Scorer (Jia et al. 2024); 3) reinforcement learning methods: Q-Insight (Li et al. 2025a), VisualQuality-R1 (Wu et al. 2025), and VQ-Insight (Zhang et al. 2025a). To ensure fairness, we retrain the IQA-based methods such as Q-Insight and VisualQuality-R1 on the LSVQ training set, using the same sampled frames as VQAThinker to adapt them for the VQA task. [^4] Since VQ-Insight is not open-sourced, we report its performance based on the official results.
@@ Conclusion
In this work, we propose VQAThinker, a reasoning-based VQA framework that leverages group relative policy optimization (GRPO) to achieve generalizable and explainable VQA. Specifically, we introduce three VQA-specific rewards: a bell-shaped regression reward for fine-grained score prediction, a pairwise ranking reward for relative quality discrimination, and a temporal consistency reward for modeling temporal degradations. Extensive experiments across ten VQA benchmarks and two quality description datasets demonstrate that VQAThinker achieves state-of-the-art performance in both in-domain and OOD settings, while also providing strong interpretability through explicit reasoning traces. For future work, we plan to scale up our reinforcement learning framework by incorporating large-scale synthetic distortion data with pseudo labels, thereby enhancing the diversity of training content and distortion types, and explore test-time scaling to further improve video quality scoring and understanding performance.
@@
YouTube-UGC (Wang, Inguva, and Adsumilli 2019) 2019 1,380 1,380 360p–4K 20 30 In-the-wild
@@ Appendix B More Details of Our Testing Benchmarks
  • KoNViD-1k (Hosu et al. 2017): KoNViD-1k is an authentic video quality assessment (VQA) dataset consisting of 1,200 unique test videos with diverse real-world distortions. All videos are sampled from the YFCC100M dataset (Flickr) (Thomee et al. 2016) based on a feature space including blur, colorfulness, contrast, spatial information, temporal information, and NIQE (Mittal, Soundararajan, and Bovik 2012). Each video is clipped from the original source content and resized to 540p with a landscape layout. The frame rates of the videos are 24, 25, or 30 fps, and the duration of each video is 8 seconds.
@@
  • YouTube-UGC (Wang, Inguva, and Adsumilli 2019): YouTube-UGC is an authentic VQA dataset consisting of 1,380 unique test videos (1,020 currently available) exhibiting a wide range of real-world distortions. The videos are sampled from YouTube using a feature space defined by spatial, color, temporal, and chunk-level variations, covering diverse content types such as HDR, screen recordings, animations, and gaming videos. The resolutions include 4K, 1080p, 720p, 480p, and 360p, with both landscape and portrait orientations. Frame rates range from 15 to 60 fps (including 15, 20, 24, 25, 30, 50, and 60), and each video lasts 20 seconds.
@@ Appendix C More Details of Our Comparing Methods
  • NIQE (Mittal, Soundararajan, and Bovik 2012): NIQE constructs a “quality aware” collection of statistical features derived from a natural scene statistics (NSS) model. These features are extracted from a corpus of undistorted natural images and fit to a multivariate Gaussian (MVG) model. The quality of a test image is then quantified as the distance between the MVG of its NSS features and the MVG fitted on natural images.
  • VIIDEO (Mittal, Saad, and Bovik 2015): VIIDEO constructs perceptually relevant space-time statistical models based on temporal frame differences and analyzes deviations from these models using NSS. The model fits the NSS features to a MVG, and the video quality is inferred from the extent of deviation.
@@
  • CLIP-IQA (Wang, Chan, and Loy 2023): CLIP-IQA leverages the intrinsic perceptual capabilities and semantic priors encoded in CLIP without requiring any training on quality labels. It formulates IQA as a vision-language matching problem through prompt engineering, where positive and negative quality prompts (e.g., “Good photo.” vs. “Bad photo.”) are constructed to evaluate visual quality. To mitigate ambiguity in natural language interpretation, a contrasting prompt pair strategy is employed to ensure robust and consistent comparisons.

### v2

@@ Introduction
Video quality assessment (VQA) (Min et al. 2024) plays a fundamental role in video processing systems, serving as the primary criterion for evaluating perceptual quality across stages such as acquisition, compression, and enhancement. Most established VQA models (e.g., VMAF (Li et al. 2018)) rely on full-reference (FR) settings, where a pristine reference video is available for comparison. In many real-world scenarios, particularly those involving user-generated content (UGC), reference videos are unavailable, making FR methods inapplicable. Therefore, no-reference (NR) VQA, which assesses quality without relying on reference videos, offers broader applicability and is the focus of this work  [^1] In this paper, we use VQA to refer to NR-VQA for simplicity..
@@
NR-VQA has witnessed significant advancements, progressing from early handcrafted feature-based methods, such as natural scene statistics (NSS) (Mittal et al. 2012a) to deep neural network (DNN) architectures, including convolutional neural networks (CNNs) (Li et al. 2019a), Transformers (Wu et al. 2022), and more recently, large multimodal models (LMMs) (Ge et al. 2024). However, most existing methods rely on supervised fine-tuning (SFT) using specific video content or distortion datasets (Ying et al. 2021; Li et al. 2019b), resulting in models that act as specialists for narrow scenarios but exhibit poor generalization to unseen video domains. To enhance generalization, recent studies have explored strategies, such as multi-dataset joint training (Li et al. 2019a), unsupervised and self-supervised learning (Cao et al. 2025), etc., aiming to improve training diversity while minimizing labeling costs. Nevertheless, these approaches often incur significantly higher training costs due to the increased data scale and model complexity.
@@
In addition, interpretability has become a crucial capability for practical NR-VQA applications. While conventional VQA models typically output singl- or multi-dimensional quality scores, real-world scenarios often require diagnostic assessments that reveal the underlying causes of quality degradation, such as compression artifacts, temporal instability, etc. Such interpretability not only enhances the reliability of VQA predictions but also enables closed-loop video processing by guiding targeted restoration strategies based on identified distortions. Recently, the rise of LMMs has advanced the development of explainable VQA, where models are typically fine-tuned on human-annotated quality instruction datasets, such as VQA2 (Jia et al. 2024) and OmniVQA-Chat-400K (Jia et al. 2025). However, these datasets are typically bootstrapped from LMM-generated descriptions and later refined by human annotators, making the final annotations susceptible to the biases and limitations of the source models. Furthermore, they often lack associated quality scores, hindering the joint learning of quality understanding and scoring, a process believed to mirror human perceptual decision-making (Mazurek et al. 2003).
@@
Developing generalizable VQA models remains a long-standing objective. Early studies primarily utilize handcrafted features based on perceptual quality priors, such as NSS (Mittal et al. 2012a), motion vectors (Konrad and Dubois 1992), optical flow (Beauchemin and Barron 1995), etc. While effective on synthetic distortions, these methods demonstrated limited efficacy on videos containing complex real-world distortions. Subsequently, DNN-based methods have been widely adopted for VQA, which typically comprise two main modules: a feature extraction module and a regression module. The feature extraction module often employs off-the-shelf networks, such as CNNs (Li et al. 2019a), Transformers (Wu et al. 2022), and LMMs (Wu et al. 2023e), which are either pre-trained on large-scale datasets (e.g., IQA datasets (Li et al. 2022)) or fine-tuned directly on the target VQA datasets (Sun et al. 2022; Wu et al. 2022) to capture the spatial and temporal quality representation. The regression module then maps the extracted features to quality scores, using either MLPs that aggregate frame-level predictions (Sun et al. 2024) or sequence modeling methods, such as GRUs (Li et al. 2019a), InceptionTime (Ying et al. 2021), or Transformers (Wu et al. 2023a), to capture temporal dependencies across frames. DNN-based methods significantly improve the performance on in-the-wild distortions, but they still suffer from limited generalization to unseen video content and distortion types (Cao et al. 2025).
To improve generalization, recent studies have explored various training strategies beyond standard supervised learning. These include mixed-dataset training with scale alignment (Li et al. 2019a), contrastive learning with proxy tasks (Chen et al. 2021; Madhusudana et al. 2023), distortion-aware reconstruction via encoder-decoder architectures (Xie et al. 2024), and learning-to-rank frameworks using pseudo-labeled video pairs (Cao et al. 2025). Although these methods improve generalization to some extent by increasing the diversity of training data, they typically incur substantial costs in constructing large-scale datasets and demand significantly higher training overhead. In this paper, we demonstrate that VQAThinker enhances generalization without increasing the number of training samples, highlighting its efficiency.
@@
Q-Insight NeurIPS’25 LSVQ (28K) 0.644 0.639 0.601 0.648 0.751 0.753 0.624 0.708 0.560 0.591 0.635 0.651
VisualQuality-R1 NeurIPS’25 LSVQ (28K) 0.795 0.796 0.716 0.744 0.784 0.792 0.732 0.781 0.717 0.730 0.765 0.776
@@
Q-Insight NeurIPS’25 LSVQ (28K) 0.310 0.326 0.372 0.384 0.256 0.268 0.218 0.206 0.547 0.564 0.334 0.340
VisualQuality-R1 NeurIPS’25 LSVQ (28K) 0.472 0.548 0.493 0.574 0.340 0.347 0.227 0.298 0.622 0.637 0.408 0.458
@@
Table 1: Performance comparison of our model against other competitive methods. The best results are highlight in boldface, the second best is underlined. NA in the table indicates unavailable results. “Overall” represents the weighted average results based on the number of videos in each dataset.
@@
Bell-shaped Regression Reward. VQA is naturally formulated as a regression task. However, existing regression-based rewards struggle to provide fine-grained training signals. Some methods use constant binary rewards (Li et al. 2025a), assigning fixed values when predictions fall within a threshold, without capturing their proximity to the ground truth. Others adopt L1-norm-based rewards (Zhang et al. 2025a), where the reward scales linearly with prediction error at a constant rate. In practice, learning becomes increasingly difficult as the predicted score approaches the ground truth. Therefore, the reward function should provide stronger and more rapidly changing signals in this fine-grained regime to better guide the model toward high-precision predictions.
@@
For video quality scoring, we evaluate performance on ten VQA benchmarks: LSVQ Test (Ying et al. 2021), LSVQ 1080p (Ying et al. 2021), KoNViD-1k (Hosu et al. 2017), LIVE-VQC (Sinno and Bovik 2018), YouTube-UGC (Wang et al. 2019), LIVE-YT-Gaming (Yu et al. 2022), CGVDS (Saha et al. 2023), LIVE-YT-HFR (Madhusudana et al. 2021), Waterloo-IVC-4K (Li et al. 2019b), and VDPVE (Gao et al. 2023). The first five datasets are considered in-domain, as they primarily consist of UGC videos with in-the-wild distortions. The remaining five datasets are treated as OOD due to their distinct content types (e.g., gaming videos) or distortion patterns (e.g., compression and enhancement artifacts). For video quality understanding, we adopt two evaluation benchmarks: FineVD (Duan et al. 2025) and Q-Bench-Video (Zhang et al. 2025b). FineVD is used to assess the model’s ability to identify quality-related distortion attributes, while Q-Bench-Video measures the accuracy of video quality descriptions.
@@
Competing Methods We compare VQAThinker with three categories of methods: 1) unsupervised methods: NIQE (Mittal et al. 2012b), VIIDEO (Mittal et al. 2015), STEM (Kancharla and Channappayya 2021), and CLIP-IQA (Wang et al. 2023); 2) supervised methods: SimpleVQA (Sun et al. 2022), FAST-VQA (Wu et al. 2022), DOVER (Wu et al. 2023b), MinimalisticVQA (IX) (Sun et al. 2024), Q-Align (Wu et al. 2023e) and VQA2-Scorer (Jia et al. 2024); 3) reinforcement learning methods: Q-Insight (Li et al. 2025a), VisualQuality-R1 (Wu et al. 2025), and VQ-Insight (Zhang et al. 2025a). To ensure fairness, we retrain the IQA-based methods such as Q-Insight and VisualQuality-R1 on the LSVQ training set, using the same sampled frames as VQAThinker to adapt them for the VQA task. [^3] Since VQ-Insight is not open-sourced, we report its performance based on the official results.
@@ Conclusion
In this work, we propose VQAThinker, a reasoning-based VQA framework that leverages GRPO to achieve generalizable and explainable VQA. Specifically, we introduce three VQA-specific rewards: a bell-shaped regression reward for fine-grained score prediction, a pairwise ranking reward for relative quality discrimination, and a temporal consistency reward for modeling temporal degradations. Extensive experiments across ten VQA benchmarks and two quality description datasets demonstrate that VQAThinker achieves state-of-the-art performance in both in-domain and OOD settings, while also providing strong interpretability through explicit reasoning traces.
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grant 62522116, Grant 62271312, Grant 62132006, and Grant 62301316, and in part by STCSM under Grant 22DZ2229005.
@@ Acknowledgements
YouTube-UGC (Wang et al. 2019) 2019 1,380 1,380 360p–4K 20 30 In-the-wild
@@ Appendix B More Details of Our Testing Benchmarks
  • KoNViD-1k (Hosu et al. 2017): KoNViD-1k is an authentic video quality assessment (VQA) dataset consisting of 1,200 unique test videos with diverse real-world distortions. All videos are sampled from the YFCC100M dataset (Flickr) (Thomee et al. 2016) based on a feature space including blur, colorfulness, contrast, spatial information, temporal information, and NIQE (Mittal et al. 2012b). Each video is clipped from the original source content and resized to 540p with a landscape layout. The frame rates of the videos are 24, 25, or 30 fps, and the duration of each video is 8 seconds.
@@
  • YouTube-UGC (Wang et al. 2019): YouTube-UGC is an authentic VQA dataset consisting of 1,380 unique test videos (1,020 currently available) exhibiting a wide range of real-world distortions. The videos are sampled from YouTube using a feature space defined by spatial, color, temporal, and chunk-level variations, covering diverse content types such as HDR, screen recordings, animations, and gaming videos. The resolutions include 4K, 1080p, 720p, 480p, and 360p, with both landscape and portrait orientations. Frame rates range from 15 to 60 fps (including 15, 20, 24, 25, 30, 50, and 60), and each video lasts 20 seconds.
@@ Appendix C More Details of Our Comparing Methods
  • NIQE (Mittal et al. 2012b): NIQE constructs a “quality aware” collection of statistical features derived from a natural scene statistics (NSS) model. These features are extracted from a corpus of undistorted natural images and fit to a multivariate Gaussian (MVG) model. The quality of a test image is then quantified as the distance between the MVG of its NSS features and the MVG fitted on natural images.
  • VIIDEO (Mittal et al. 2015): VIIDEO constructs perceptually relevant space-time statistical models based on temporal frame differences and analyzes deviations from these models using NSS. The model fits the NSS features to a MVG, and the video quality is inferred from the extent of deviation.
@@
  • CLIP-IQA (Wang et al. 2023): CLIP-IQA leverages the intrinsic perceptual capabilities and semantic priors encoded in CLIP without requiring any training on quality labels. It formulates IQA as a vision-language matching problem through prompt engineering, where positive and negative quality prompts (e.g., “Good photo.” vs. “Bad photo.”) are constructed to evaluate visual quality. To mitigate ambiguity in natural language interpretation, a contrasting prompt pair strategy is employed to ensure robust and consistent comparisons.